OpenStudy (anonymous):

as x goes to infinite what is the limit of (e^x + x ) ^1/x

OpenStudy (rock_mit182):

this is a tough one atleast for me

OpenStudy (freckles):

did you try this: \[\text{ \let } y=(e^x+x)^\frac{1}{x}\] \[\ln(y)=\frac{1}{x} \ln(e^x+x)\] so \[y=e^{\ln(y)}=e^{\frac{1}{x}\ln(e^x+x)} \] Then just to find \[\lim_{x \rightarrow \infty }y=e^{\lim_{x \rightarrow \infty} \frac{\ln(e^x+x)}{x}}\]

OpenStudy (freckles):

then use l'hospital a couple of times